brandybriskey1:

a train needs 4 hours to go from the first station to the second. if the third station is 120 miles away and the total travel time must be no more than 7 hours, write an inequality that would let you find the slowest average speed, v, the train could have.

Vocaloid:

so let's write a sum for the total travel time and set that sum less than or equal to 7 for the first-second segment, it's just t_1 = 4(given) for the second-third segment, then distance = rate*time, so 120 = v*t_2 and t_2 = 120/v let \[t_1 + t_2 \le 7\] and solve for v

Eiwoh2:

:P Welp. Can't you close all these unanswered questions?

Vocaloid:

Closed due to answered.